### v1

# LaCour!: Enabling Research on Argumentation in Hearings of the European Court of Human Rights

Abstract
Why does an argument end up in the final court decision? Was it deliberated or questioned during the oral hearings? Was there something in the hearings that triggered a particular judge to write a dissenting opinion? Despite the availability of the final judgments of the European Court of Human Rights (ECHR), none of these legal research questions can currently be answered as the ECHR’s multilingual oral hearings are not transcribed, structured, or speaker-attributed. We address this fundamental gap by presenting LaCour!, the first corpus of textual oral arguments of the ECHR, consisting of 154 full hearings (2.1 million tokens from over 267 hours of video footage) in English, French, and other court languages, each linked to the corresponding final judgment documents. In addition to the transcribed and partially manually corrected text from the video, we provide sentence-level timestamps and manually annotated role and language labels. We also showcase LaCour! in a set of preliminary experiments that explore the interplay between questions and dissenting opinions. Apart from the use cases in legal NLP, we hope that law students or other interested parties will also use LaCour! as a learning resource, as it is freely available in various formats at https://huggingface.co/datasets/TrustHLT/LaCour.
keywords:
ECHR, Legal arguments, Hearings, Corpus
1 Introduction
What can we learn about law and legal argumentation from court judgments alone? Contemporary research addresses empirical legal questions (e.g., which arguments are used) or legal NLP questions (e.g., predicting case outcomes) by relying on the availability of the final ‘products’ of each case, the court decisions (Habernal et al, 2023; Medvedeva et al, 2020). The European Court of Human Rights (ECHR) is a prominent data source, as its decisions are freely available in a large amount, along with the metadata of the violated articles and other attributes. This makes ECHR a popular choice among NLP researchers (Aletras et al, 2016; Chalkidis et al, 2020). However, whether or not the legal arguments in ECHR’s cases are created as a part of legal deliberation or are created post-hoc after reaching a decision remains an open (and partly controversial) question.
In order to better understand the legal argument mechanics, that is which arguments of the parties were presented, discussed, or questioned, and thus might have influenced the case outcome, we must take the oral hearings into account. We witness that the availability of oral hearing transcripts of the U.S. Supreme Court enables further legal research (Ashley et al, 2007). However, empirical research into the interplay of arguments at the court hearings and the final judgments has been so far impossible for the ECHR, as there are no hearing transcripts available.
This paper addresses this research gap. Given only the raw video material recorded during selected ECHR hearings, we ask to which extent can we semi-automatically create a high-quality multi-lingual corpus that would enable further empirical legal NLP research. The challenges that we need to tackle include second-language speakers in legal jargon, missing speaker attribution (which judge or party presents their arguments/questions and when), low-quality court videos with technical glitches, and no alignment between the hearings and the final published court decisions. We address these challenges by combining automatic speech recognition with extensive manual corrections and annotations and create transcripts that are structured, as accurate as possible and, last but not least, easy to read.
We present LaCour!, the first corpus of partly manually-corrected and speaker-aligned multi-lingual textual transcripts of hearings of the ECHR. To showcase an exemplary scenario in which this corpus can be used, we study the correlation between the interactive part of a hearing and the formation of a judicial opinion in the judgment. More specifically, we focus on the connections between the questions posed by the judges to the parties during the hearings and the possible judicial opinions of these judges later on. We would like to know explicitly whether it is possible to see a foreshadowing of this opinion already during the hearing.
We hope that LaCour! will open up new research possibilities for the NLP community to look at the whole process of a case. In addition the transcripts can also benefit the empirical legal community, which may wish to study specific hearings in written form along with the corresponding judgments.
2 Related work
As legal NLP is a large subfield of NLP covering a wide range of topics, here we focus mainly on works dealing with hearing transcripts. We cover both the U.S. Supreme Court as well as the ECHR.
2.1 Hearings of the U.S. Supreme Court
For the U.S. Supreme Court there are two major resources providing the public with material on oral hearings. The first one is the Oyez [^1] https://www.oyez.org/ website, which is the brainchild of Jerry Goldman (Goldman, 1998). The website contains and curates oral hearings before the U.S. Supreme Court as audio and textual versions and includes entries that date back to 1955. Since its original creation, the project has continued as a collaboration between Cornell’s Legal Information Institute, Justia and Chicago-Kent College of Law. The Oyez Team has taken great effort to transfer material from the pre-digitial era into a digital and publicly available format. Most of the source material has been curated and enriched with additional information, like audio-aligned transcripts, case summaries and Supreme Court opinions. Goldman (1998, p. 34) highlights that it is crucial to have the possibility to actually listen to a hearing itself, because it conveys more meaning and nuances than only a transcript would, as these recordings “are rich in emotive content, which may have a material bearing on the attitudes and directions in a justice’s line of argument.” The task of transcription was delegated to professional companies. [^2] https://www.oyez.org/about cites three companies, Heritage Reporting, Tech Synergy and Alderson Reporting for contributing transcriptions All recordings from 1955 onwards have been digitized, resulting in a total of more than 14,000 hours of audio material and over 66 million words. [^3] https://www.oyez.org/history The database was intended as a learning and researching resource for students as well as legal scholars.
The second important resource for material concerning the U.S. Supreme Court hearings is courtlistener [^4] https://www.courtlistener.com/ which is part of the Free Law Project. [^5] https://free.law/about The site provides access to legal opinions, various documents and oral argument recordings from federal, state and specialty courts. Although oral arguments are available in digital format, no transcripts are included. There are also other websites curating legal data for the public, like the Supreme Court Database (Spaeth et al, 2022) or the Caselaw Access Project, [^6] https://case.law/ however, they do not include oral arguments in their provided material.
Easily accessible and extensive collections of all court material naturally play an important role for legal scholars. But the transcripts alone are also of great importance for empirical research, as they enable such research for NLP in the first place. Some approaches have used the transcripts of the oral arguments to gain insights into the argument structure for the purpose of tutoring (Ashley et al, 2007), others have tried stance detection (Bergam et al, 2022) or machine learning-based approaches to study the judges’ attitudes (Dickinson, 2018).
2.2 Hearings of the ECHR
On the European side of legal analysis in the context of NLP, a popular choice of investigation is the ECHR. [^7] https://www.echr.coe.int/ This court is a frequently used source for analysis, not only because of the availability of its documents, but also because of its importance and reputation as a supranational body. It has been in existence since 1953 and deals with applications brought by individuals or groups, in rare cases even states, against states that are contractors to the European Convention on Human Rights (“The Convention”). The Convention contains several articles [^8] https://www.echr.coe.int/european-convention-on-human-rights concerning different aspects of human rights and applications can refer to multiple of those articles. The court receives over 50,000 applications every year which are proceeded according to importance and urgency. For roughly 30 Chamber or Grand Chamber cases each year, the court holds a public hearing which is recorded.999more information on how an application is processed can be found in https://www.echr.coe.int/documents/d/echr/your_application_eng
In contrast to the good availability of the U.S. Supreme Court argument transcripts, no such data is available for the ECHR. The ECHR solely publishes judgments or decision through its HUDOC database. [^10] https://hudoc.echr.coe.int/ For this reason, most works pose research questions that can leverage information available in the HUDOC data collection.
A frequently pursued line of research examines the argumentative nature of the judgments (Mochales-Palau and Moens, 2007; Mochales and Ieven, 2009; Mochales and Moens, 2011; Poudyal et al, 2020); we refer to a recent work by Habernal et al (2023) for a survey of this research area.
Another popular use of ECHR material is Legal Judgment Prediction (LJP), most notably investigated by Aletras et al (2016). Besides classifying cases in a binary manner for violation of a specific Article, they also introduce a publicly available corpus with 584 cases. All cases in this dataset are in English, relate to Articles 3, 6 and 8 of the Convention and are balanced in regards to violation and non-violation. Chalkidis et al (2019) collect 11.5k cases in a similar manner and use them for binary classification, multi-label classification and case importance detection with the help of neural models. Medvedeva et al (2020) use a similarly sized portion of ECHR cases and experiment with only using information available before the final judgment to simulate actual prediction rather than classification. In an additional experiment, they also investigate a connection of a specific judge being part of the Chamber and the decision of the case. The latter experiments even achieved a surprisingly high correctness for certain Articles of the Convention. However, prediction for ECHR cases still suffers from the issue that argumentative text is only available after the judgment has been passed and anything using the judgment documents is already biased towards the outcome. Moreover, as Villata et al (2022) pointed out in their review of (Medvedeva et al, 2020), classifying legal cases is a different and more complex matter than a simple binary classification.
Finally, ECHR documents are also a popular source used in large dataset collections or benchmarks (Chalkidis et al, 2022; Henderson et al, 2022), which made them a standard and go-to option for many works in legal NLP. And due to the amount of available text, they have also been used as training material for domain specific language models, like LEGAL-BERT (Chalkidis et al, 2020) which have opened up new possibilities for handling legal texts.
However, both legal argument mining and legal judgment prediction focus on the written documents provided by the ECHR through HUDOC. Because there are no transcripts of the public oral arguments in hearings in the ECHR available, to the best of our knowledge, no such research has been conducted in this area and although the hearings are publicly available, they have not been scrutinized in NLP research to date. The hearings also present the gap in the current representation of legal proceedings in the ECHR. Although held only on a proportionally small number of cases, oral hearings can provide insight into the process that takes place between an application and the final judgment.
Figure 1: An illustration of a typical hearing structure. Variations in the structure are possible, e.g. occasionally the announcement is missing or no questions are asked
3 Creating the LaCour! corpus
A typical hearing before the European Court of Human Rights is structured into several phases which are outlined in Figure 1. The judge president opens the hearing, introduces basic information concerning the case and acts as a moderator throughout the session. They also announce the predetermined order of addresses and call the first party. Their job is also to remind the speakers when their allocated time has passed. After each party has taken their turn to address the Court and present their interpretation, the judges are called to ask questions to the parties. In earlier hearings, the judge president would then ask the parties to immediately reply to those questions as well as to the other party’s remarks, but in more recent hearings, a short break is usually permitted to allow for some preparation. In the most recent hearings there is even an opportunity for the judges to ask follow-up questions again after the replies to the first round of questions. After these replies, the judge president delivers the closing statement. This debate-like, moderated structure gives the hearings a ceremonial character. But nevertheless, the language used is different, as it is not as formal as that of written legal texts, even though “Legalese” is still present in the vocabulary used.
3.1 Hearing videos
As our source material, we collected all videos of hearings - called webcasts - that are available on the website [^11] https://www.echr.coe.int/webcasts-of-hearings. At the time we started the collection, the available hearings ranged from 2007 to 2022 with 3 different audio tracks: the original version without interpretation, an English interpretation and a French interpretation. There are no official transcripts of the hearings and the videos do not contain any subtitles. The only available material are live recorded videos with the live recorded interpretations as an optional audio track.
A single hearing is not necessarily held in a single language. Besides the official languages of the Court, English and French, participants are often allowed to speak in their native language. Consequently, a certain percentage of hearings (around 20% in our subset) are not held in one or two, but several spoken languages. Both the English and French interpretation audio tracks are recordings of the live interpretation which were created at the time of the hearing and do not sound as fluent as the original. Although we have no way of knowing which version a participant or observer heard at the live hearing, we chose to focus on the original version, which contains no interpretation or translation. Our initial collection contained 289 hearings between 2007 and 2022 with the original audio in various languages at the time we scraped the website.
We probed the quality of the videos and noticed that most videos before 2012 are perceived as worse than recent ones and prone to technical errors, like sound dropouts, microphone failures and incorrect audio tracks. A preliminary transcription test with some of the hearings of the early years also revealed a poor quality. In addition, there are several hearings with parts of the hearing missing or cut off. Therefore, we decided to only use hearings starting from the year 2012.
Figure 2: Illustration of the transcript generation pipeline. Yellow elements indicate an automatic processing, while green elements indicate manual annotation and correction. The associated code is available on https://github.com/trusthlt/lacour-generation
3.2 Generating transcripts
The very first step in the processing pipeline, illustrated in Figure 2, is the speaker diarization, which is applied to obtain a schedule of who is speaking at which time. The diarization also takes care of overlapping speech that occurs in rare cases. For this task, we ran pyannote [^12] https://huggingface.co/pyannote/speaker-diarization (Bredin et al, 2020) which returns a fine-grained schedule of speech segments and a supposed speaker identification number. We clustered the output of pyannote by speaker change into larger chunks to create a “speaker schedule” that resembles the order of speeches in the hearing.
To account for the different languages spoken, we originally wanted to use automatic language detection on those segments. However, initial trials with several wav2vec2 models to detect the spoken language did not yield sufficiently reliable accuracy. And even though the field of speech and language recognition has seen large improvements with models like Whisper (Radford et al, 2023), we still encountered a number of issues with the language detection in Whisper as well. One major problem is the fact that participants of the hearings, especially the judges, tend to speak both English and French and not seldom use both languages in a single sentence. A common example of this is “Merci beaucoup, Madame. Judge [Name], please.” The automatic language detection for utterances like this failed frequently. The second most common problem is related to the language proficiency of the speaker. With English and French being the official languages, many speakers before the Court do not speak in their native language. And although recognizing the spoken language in these cases might not be an issue for a human, the automatic language detection often picks up the native language of the speaker, rather than the actual spoken language. We suspect the detection model might be biased towards native speakers.
To circumvent this unreliable detection, we manually identified the languages for each section in our coarse-grained speaker schedule. We hired two research assistants to listen to parts of the speaker segments in order to identify the spoken language as well as the affiliation of the speaker, which would later be needed to distinguish a party’s address from the judges’ questions. Any conflicting identification by the annotators was discussed until a correct tag was agreed on. Additionally, this manual step allowed us to correct small diarization errors of the framework. The resulting speaker schedules contain segments which are marked with timestamps, as well as the spoken language (ISO 639-1) and the role of the speaker. For the roles, we decided to distinguish between the parties appearing in court. These are Announcer, Applicant, Government, Judge President (“JudgeP”), Judge and Third Party. If there are multiple parties as for example multiple applicants in a case, the tags do not distinguish between them. In the case “Savickis and Others v. Latvia (no. 49270/11)” for example, the representatives for “Savickis” and the representatives for all other applicants would be tagged simply as “Applicant”.
The central component of the pipeline is the transcription itself. Our method for generating a transcript involves using Whisper (Radford et al, 2023) on a single entry of a speaker schedule, passing the previously identified language as a parameter instead of using the model’s language detection. For long segments of speech this produces a single large block of text. To ensure better alignment between timestamps and transcript, we switched to using a sentence segmenting version of Whisper [^13] https://github.com/jianfch/stable-ts to generate more accurately aligned sentence segments of the transcript. The resulting transcript is divided into speaker segments which are additionally divided into fine-grained sentence segments. One additional round of manual correction was necessary to properly separate instances of speaker overlap, as the model generates inaccurate transcripts when two people are speaking simultaneously. Segments with potential overlap were tagged automatically and evaluated by our student research assistants to resolve these conflicts.
Figure 3: Excerpt of the transcript to webcast of Savickis and Others v. Latvia (no. 49270/11), webcast id 4927011_26052021
Even though this transcript version is sufficient for understanding the hearings, we decided to further refine it by adding the names of the judges and correcting the transcript. To obtain the names, we collected lists of hearing participants from linked documents in the HUDOC databse titled “Forthcoming Hearing” and verified that the participants listed in them also appear in the “Judgement” document. This step was necessary to spell the names of the representatives correctly, particularly when they are mentioned in the transcript. The names of the judges could be used later to link judges’ opinions in the final judgments with questions asked by them. Although the produced transcripts are surprisingly accurate, legal terminology and named entities are frequently transcribed incorrectly, e.g. special terminology such as “tax arrears” is often transcribed as “tax errors”.
We hired two student research assistants proficient in English and French and tasked them with manually correcting the parts spoken by the judges and the judge president. They reported an average effort of 1,5 hours per hearing and corrected the parts of the entire corpus in approximately 300 hours. Due to limited resources, speeches by other speaker roles (“Applicant”, “Government”, “Third Party”) and in other languages remain uncorrected. Our completed transcript corpus comprises tagged speech segments with tagged sentence-aligned transcriptions, which are manually corrected for the labels “Judge” and “Judge President”, but uncorrected for all other labels. Figure 3 displays an example of a transcript. Obtaining the final partially corrected transcript from the video version required the work of 4 student research assistants and took about 10 months. We call our new corpus LaCour! in honor of the catchy announcement at the beginning of each hearing.
3.3 Linked documents
In order to represent a large part of the entire process that surrounds a hearing and case as adequately as possible, the hearing needs to be considered in its context. We therefore also collected the most important documents that accompany a hearing as a supplementary text resource. Although documents relating to written communication prior to a hearing are scarce, there is almost always a concluding document containing the “Judgement” by either the Grand Chamber or the Chamber. Since those documents summarize the entire proceedings surrounding the case at hand, they can play an important role for studying the argumentation and course of events. In addition, they detail the majority opinion of the court and the decision, as well as dissenting or concurring opinions by individual judges.
We collected documents from the HUDOC database that are associated with any application number that was attached to a hearing webcast. Due to different document types, there are multiple documents associated with a hearing. Some manual adjustments to the matching were required, because the official database occasionally links to an incorrect document. We exclusively retrieved English documents, but it is also possible to find the link to a French document, should it be available. During the process of acquiring the matching judgment files, we noticed some of the applications were deemed inadmissible or were withdrawn by the applicant after the hearing. Because having a hearing transcript without a conclusion in form of a judgment appeared insufficient, we excluded all hearings without a judgment file. The list of excluded hearings can be seen in Appendix A along with the reason for exclusion.
All judgment files are provided with their respective case-id and application number, the latter of which serves as a link to the hearing and webcast-id. A webcast only has a single id, while there can be more than one application number connected to a hearing or a case. It should also be noted that there is overlap [^14] We found an overlap based on the item-id for 190 out of 469 documents of these documents to the corpus by Chalkidis et al (2019), but unfortunately there is no overlap to the documents annotated by Habernal et al (2023). This is primarily because LaCour! mostly contains documents with an importance score of 1, meaning “key cases” and only from a limited date range. The case-ids and application numbers can also serve as a key to identify and collect more material, because it is the same identification as used by the court.
Additional quality control and limitations of current transcripts in LaCour!
Our transcriptions are not meant to reflect verbatim what is said in the audio, but to aid the understanding of the hearing in text form. We therefore do not correct the text to reflect stuttering, sentences that have been slightly corrected during speaking, filler words or incorrect grammar. But because Whisper relies on an internal language model, the generated sentences are sometimes grammatically correct, even though the actual spoken sentence is not. This is especially true for the parts we did not manually correct, which concerns everything spoken by the roles “Applicant”, “Government” and “Third Party”. Manually corrected and modified parts are sections spoken by “Judge” and “Judge President”, as well as all language tags. The remainder of the text has been left untouched and is a direct result of Whisper. Several format and completeness checks were additionally run on the files to ascertain that no tags are incomplete or missing and the judges’ names match the ones listed as attendants.
Dataset accessibility for NLP and explorative studies
In order to enable easy dissemination and also provide an easily readable version of a transcript, we provide LaCour! in different file formats. For comfortable reading, we refer to the concatenated version [^15] https://github.com/trusthlt/lacour-corpus or our online preview page. [^16] https://www.trusthlt.org/lacour/ The sentence aligned version is available in xml and text format on huggingface. [^17] https://huggingface.co/datasets/TrustHLT/LaCour At this point we would like to emphasize that the transcripts are unofficial interpretations of the hearing webcasts. The webcasts in conjunction with the provided documents are the only official presentation of the proceedings.
3.4 LaCour! statistics
The final corpus of hearing transcripts contains 154 hearings ranging from 2012 to 2021 transcribed in their original version and linked to the respective final judgment documents. The total duration of the video content is 267 hours, 20 minutes and 33 seconds with a file size of 153 GB. All 154 transcripts together contain 2,161,833 word tokens. Based on duration of the speeches, a large portion of the transcripts are in English, with French as the second most frequent language. The distributions are illustrated in Figure 4 with the duration of other languages detailed in Figure 4.
On average, each hearing contains 14,038 word tokens and has an average sentence count of 557. With a standard deviation of 3,519 tokens, the majority of the transcripts are roughly of equal length. Exact statistics can also be found in Appendix A. Looking at the amount of tokens per hearing categorized by party outlined in Figure 4, one can see that there are several hearings with a large than average amount of tokens and very few hearing with a comparably smaller amount of tokens. If a third party is involved in the hearing, their part will usually be short. For applicant and government, the amount is also roughly equal. A closer look at the balance between applicant and government is also illustrated in Figure 4, which shows the absolute difference between the percentage of tokens spoken by the two opposing parties based on the length of the entire hearing. Apart from a few outliers, 85% of all hearings remain within a maximum of 20% difference between the two opposing parties. This means that generally, hearings have a balanced amount of time allocated for each party. The outliers can mostly be traced to the presence of more than one party on either side, which means that either there were multiple applicants or multiple respondents involved with fairly allocated speech time.
(a) Share of main languages based on time
(b) Duration of other languages in the share “Other”
(c) Distribution of the amount of tokens of each hearing grouped by role label
(a) Share of main languages based on time
(b) Duration of other languages in the share “Other”
(c) Distribution of the amount of tokens of each hearing grouped by role label
(d) Dominance one party has over the other (absolute) based on the share of tokens
Figure 4: Distribution of the languages and roles in the corpus based on time (a,b) and amount of tokens (c,d)
4 Preliminary experiments
In order to test and showcase our new corpus, we explore potential tasks which emerge with the new type of data for ECHR court proceedings. The LaCour! corpus partly bridges the gap between an application and a judgment, so we would like to investigate the impact of the hearing on the final outcome. While annotating the transcripts, we observed that some judges ask questions in almost every hearing, while others remain quiet. The questions often ask for some clarification on specific issues. The importance of questions in legal proceedings has also been studied by Johnson et al (2009) in the context of the U.S. Supreme Court. Their work, alongside our subjective observations during the annotation process, inspired us to take a closer look at the question-asking part of the hearings. Together with the questions asked during the hearings, the opinions give direct insight into an individual judge’s view on the case. We want to investigate whether there is a connection between a judge having a question during the hearing and a judge having a dissenting or concurring opinion in the subsequent judgment.
Judgments are formed by the Chambers and usually the majority of judges in the current Chamber have to agree on the verdict to form a majority opinion. There are multiple votes depending on the articles of the convention and aspects which are discussed. If an individual judge disagrees with the majority on a topic, they are asked to append their opinion in written form to the final judgment document. The most common types of judicial opinions are “dissenting”, which occur when a judge disagrees with the majority opinion and “concurring”, when a judge agrees with the outcome but for different reasons. In LaCour!, the linked judgment documents very frequently include these opinions. The most common type of opinion is “DISSENTING”, followed by “CONCURRING”. Some opinions are simply declared as “opinion” and some are partly dissenting and partly concurring. An opinion can be formulated by one or multiple judges. We extracted opinions of our linked documents with the help of regular expressions and further tagged them as “DISSENTING”, “CONCURRING”, “PARTLY” or simply “OPINION” if no further details are given. We excluded documents exclusively in French and opinions by judges who were not listed as present in the hearing and collected 758 opinions, 360 of which are unique. The distribution of the labels is shown in Table 1.
label unique opinions judge with opinion
DISSENTING 213 507
CONCURRING 94 182
OPINION 44 60
PARTLY 9 9
total 360 758
Table 1: Opinions with labels sourced from the judgment documents associated with the hearings; an opinion can be formulated by multiple judges
First, we conducted the exact Fisher test on all participating judges of our 154 hearings which are also listed in the judgment. We extracted matching questions from our transcripts by taking all content tagged with “Judge” [^18] This approach ignores potential questions asked by the Judge President, but it allowed us to extract questions without additional annotation or manual work and filtering questions with less than 25 tokens. This resulted in 651 questions (522 of them in English) with an average of 248 tokens. A cross tabulation of the extracted material, limited to only participants listed in both hearing and judgment, resulted in the numbers listed in Table 2.
has opinion
has question False True total
False 1304 530 1834
True 361 228 589
total 1665 758 2423
Table 2: Cross tabulation of participating judges in a hearing which do or do not have a questions in the hearing and an opinion in the associated judgment document
[…]
Based on these observations we inspected the connection between the content of question and opinion for the instances in which there was both a question in the hearing and an opinion in the judgment. We extracted this subset and limited it to only English questions, leaving us with 193 pairs. We then created sentence embeddings of each question and each opinion with a pre-trained Sentence-BERT (Reimers and Gurevych, 2019) model. [^19] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 The task of matching the questions to their opinions is formulated as a ranking task by finding the opinion embeddings with the highest cosine similarity for each question embedding. The pool of potential matches is limited to the amount of opinions we sourced (193). The precision@1 for this task is 0.16 and this result suggests that the content of question and opinion is not unrelated and should be investigated further.
As a more complicated task, we wanted to examine if it is possible to classify the type of the opinion based only on the question. A simple logistic regression or a multilayer perceptron (MLP) using the sentence embeddings as features was unable to learn anything from the set of questions labeled with the opinion tags shown in Table 1. In fact, we were unable to learn a classifier based on whether a question is followed-up with an opinion or not. However, since the opinion ranking task seems to suggest a connection on the basis of the sentence embedding, we further train a pre-trained Sentence-BERT model using contrastive loss and pairs of questions as an input. Although we were also unable to achieve results that outperform the majority baseline, we provide details for all our experimental setups in Appendix B.
Analysis and discussion
Our experimental results are preliminary to some extent. First, we used simple tools and models which can run with very low resource requirements. It is certainly possible that larger models are able to detect hints towards the ensuing opinion in the question. Second, our experiments do not consider the context of the questions, so future work might look into ingesting the remainder of the hearing or additional information. Overall, we show that our new LaCour! corpus allows for new kinds of interesting investigations that combine hearings with data from judgments.
5 Conclusion
We introduced our new corpus LaCour! which considers oral court proceedings in the ECHR for the first time. This corpus is available to the research community and enables the empirical analysis of ECHR court hearings. In addition to the transcribed and partly manually corrected text from the audio, we provided sentence-level timestamps and manually annotated role and language labels. Our collected data aims to reduce the gap that exists when considering ECHR proceedings as a whole, because it provides insights into the processes that takes place between the submission of an application and the final decision. The dataset characteristics can also enrich existing datasets or enable the formation of new datasets, which we showcased with our subset of questions and opinions. Although our preliminary experiments were limited to simple statistical tests and general sentence-embedding models, there are indications that there is possibly a connection between the questions asked in a hearing and opinions on a judgment. We believe that these experiments could be expanded on and our findings warrant further investigation. Apart from the use cases in legal NLP, we hope that law students or other interested parties will also use our data as a learning resource.
Limitations
There are some general limitations which originate from the material we worked with. One such limitation is that even though we try to illustrate most of the process of an application and case, we chose to use the original audio presented at the associated hearing. Although in our opinion, this is the best choice for understanding the speeches by applicants who did not speak English (or French), it should be noted that the judges and attendants of the hearings do not necessarily listen to the original audio. There is no way of knowing which interpretation a judge has heard, it could be either French or English. It is therefore not possible to draw conclusions from the choice of words used in the speeches and the judgments or opinions given by the judges of the Court. All names, institutions and parties were taken from the official documents provided in the HUDOC database. The only sections which are tagged with a name are those spoken by either “Judge” or “Judge President”. This enables linking questions to opinions in the judgment documents. However, these tags should not be used to automatically profile individual judges. We would also like to repeat that a large part of the transcript is the uncorrected output of a Whisper model. It is not unlikely that named entities and legal vocabulary are incorrectly transcribed. The quality of the transcriptions has not been evaluated for languages other than English and French. It should also be noted that under no circumstances do we wish to replicate or emulate a judge. We strongly believe that judgments concerning the lives of people should also be at the hand of an actual person.
Ethics statement
This research has been approved by the ethical review board of the Technical University of Darmstadt.
Acknowledgements
The independent research group TrustHLT is supported by the Hessian Ministry of Higher Education, Research, Science and the Arts. This work has been funded by the German Research Foundation as part of the ECALP project (HA 8018/2-1) and by the hessian.AI ConnectCom grant. We would also like to thank our research assistants who helped with the creation of the LaCour! corpus: Laura Boyette, Yassine Thlija, Marwa Naibkhil and Leonard Bongard.
    Appendix B Details on preliminary experiments
    The following training setups were considered for our preliminary experiments. The full dataset of questions and opinions is also available on https://github.com/trusthlt/lacour-questions-opinions and contains all 2423 pairs of questions and opinions (or lack thereof) with a label for the opinion type and a boolean value for the presence of a question and an opinion.
    The first setup (Table 5) considers all questions and uses logistic regression to classify a) whether there was an opinion or not and b) the type of opinion. The questions are transformed into sentence embeddings with Sentence-BERT (all-MiniLM-L6v2).
    model LR MLP
    train size 304 304
    dev size 76 76
    test size 95 95
    n layers 1 5
    input size 384 384
    batch size {64, 32, 16} {64, 32, 16}
    optimizer Adam Adam
[…]
    random seed 5113 5113
    epochs 50 50
    random baseline (accuracy) [%] 43.86 43.86
    random baseline ($F_{1}$ score) [%] 35.15 35.15
    majority (all 0) baseline (accuracy) [%] 59.65 59.65
    majority (all 0) baseline ($F_{1}$ score) [%] 0.00 0.00
    model best (accuracy) [%] 59.65 59.65
    model best ($F_{1}$ score) [%] 0.00 0.00
    Table 5: Experimental setup and training parameters for logistic regression (LR) and multilayer perceptron (MLP)
    Due to the simple setup of these experiments, a negative result was expected. These experiments were also conducted to see if there is an artifact or leakage of the label in the question.
    The second setup (Table 6) performs further training on a pre-trained sentence transformer. In this case, the pre-trained model ‘all-MiniLM-L6-v2’ is used as a base model and trained in a contrastive manner. A sentence transformer ingests a pair of two sentences and outputs their cosine similarity. The idea is that similar sentences have a high cosine similarity. In our case, the training objective is that pairs of questions which are followed up by an opinion should be closer to each other, while question pairs with unequal labels should be further away from each other. We used a simple contrastive loss for this and constructed similar pairs based on whether the question appears in the same hearing and carries the same binary opinion label (whether a question was followed up with an opinion or not). Negative examples are question pairs which appear in the same hearing but do not have the same opinion label. Additional ‘soft’ samples could be generated by also considering questions from different hearings with equal or unequal labels. The trained model was evaluated on a test set with a binary classification evaluator which only considers unseen hard negative and positive samples.
    sample generation ‘hard’ samples ‘soft’ samples
    train size 942 84302
    evaluation size 182 182
    loss contrastive constrastive
    tokenizer roberta-base roberta-base
    batch size 64 64
    random seed 5113 5113
    epochs 50 10
    random baseline (accuracy) [%] 43.41 43.41
    random baseline ($F_{1}$ score) [%] 56.44 56.44
    majority baseline (accuracy) [%] 61.54 61.54
    majority baseline ($F_{1}$ score) [%] 76.19 76.19
    model before training (accuracy) [%] 62.64 62.64
    model before training ($F_{1}$ score) [%] 76.71 76.71
    model best (accuracy) [%] 61.54 60.99
    model best ($F_{1}$ score) [%] 75.86 75.77
    Table 6: Experimental setup for training a sentence transformer model initialized from all-MiniLM-L6-v2 with contrastive loss using question pairs as hard and soft samples
    html>